prove the following result
tan^-1(2/3)=1/2 tan^-1(12/5)

Grade:12

1 Answers

Rinkoo Gupta
9 years ago

tan-1(2/3)=(1/2) .2tan-1(2/3)

=1/2. tan-1(4/3)/(1-4/9)

=(/12).tan-1(4/3)/(5/9)

=1/2tan-1(12/5)

Thanks & Regards

Rinkoo Gupta

AskIITians Faculty